# Enhanced DACER Algorithm with High Diffusion Efficiency

### v1

@@ 3.1 Q-Gradient Field Guided Denoising
Proof. See Appendix A.
Consequently, Theorem 1 inspires a novel design for SDE-based policy functions. By focusing on maximizing the expected Q-value, these policies can effectively learn a soft policy distribution.
Meanwhile, langevin dynamics provides an efficient method for sampling actions from Boltzmann distributions hinton2002training :
$\pi(\boldsymbol{a}|\boldsymbol{s})=\frac{e^{\frac{1}{\alpha}Q(\boldsymbol{s},\boldsymbol{a})}}{Z(\boldsymbol{s})},$ (10)
where $\alpha>0$ is temperature coefficient, $Q(\boldsymbol{s},\boldsymbol{a})$ is state-action value function and $Z(\boldsymbol{s})$ is the partition function that normalizes the distribution. The formula derived by taking the partial derivative of both sides of Eq. (10) with respect to $\boldsymbol{a}$ can be expressed as
$\nabla_{\boldsymbol{a}}\log\pi(\boldsymbol{a}|\boldsymbol{s})=\frac{1}{\alpha}\nabla_{\boldsymbol{a}}Q(\boldsymbol{s},\boldsymbol{a}).$ (11)
Substituting Eq. (11) into Eq. (7), we can obtain the sampling process for $\pi(\boldsymbol{a}|\boldsymbol{s})$:
$\boldsymbol{a}_{t-1}=\boldsymbol{a}_{t}+\frac{\delta_{t}}{2\alpha}\nabla_{\boldsymbol{a}}Q(\boldsymbol{s},\boldsymbol{a})+\sqrt{\delta}_{t}\boldsymbol{\epsilon}.$ (12)
Equation (12) can be regarded as a trivial solution to the SDE-based policy that satisfies Theorem 1. This connection suggests using $\nabla_{\boldsymbol{a}}Q(\boldsymbol{s},\boldsymbol{a})$ as an auxiliary learning objective to guide the training of the SDE-based policy, helping to ensure the policy avoids sharp local optima and maintains representational diversity to prevent mode collapse. The reason we do not rely solely on $\nabla_{\boldsymbol{a}}Q(\boldsymbol{s},\boldsymbol{a})$ to train the score function is that training based exclusively on this gradient can lead to instability. Even minor noise in the $Q$-values may prevent the algorithm from converging to the optimal policy ding2024diffusion . Therefore, we use it only as an auxiliary training objective.
3.2 Time-weighted $\nabla_{\boldsymbol{a}}Q(\boldsymbol{s},\boldsymbol{a})$
In the reverse-time SDE framework, directly maximizing the expected Q-value yields an entire family of valid solutions. However, the lack of strong intermediate supervision can lead to suboptimal solutions, such as local optima or mode collapse. To address this, we incorporate the gradient term $\nabla_{\boldsymbol{a}_{t}}Q(\boldsymbol{s},\boldsymbol{a}_{t})$—which corresponds to Langevin dynamics, the most trivial member of this solution family—as an auxiliary objective when training the score network. However, the main issue is that $\nabla_{\boldsymbol{a}_{t}}Q(\boldsymbol{s},\boldsymbol{a}_{t})$ is independent of the diffusion time step, whereas the score function is not. This time invariance cannot meet the denoising requirements that change throughout the diffusion process. Specifically, in the later stages of diffusion, the denoising intensity should naturally decrease as the action distribution approaches the optimal policy.
To address this issue, we introduce a time-dependent weight that modulates the influence of Q-gradient guidance based on the diffusion time step, allowing for more precise control over the denoising process. Inspired by the design approach for the step size $\delta_{t}$ in Eq. (7), we can similarly design $w(t)$ using the commonly employed exponential decay function welling2011bayesian ; teh2016consistency :
$w(t)=\exp(c\cdot t+d),$ (13)
where $t$ denotes the current diffusion step; following common practice, we keep the hyperparameters $c$ and $d$ fixed at 0.4 and -1.8 across all experiments.
Furthermore, during the training process, the instability in the estimation of the Q function leads to fluctuations in the numerical value of $\nabla_{\boldsymbol{a}_{t}}Q(\boldsymbol{s},\boldsymbol{a}_{t})$, which in turn affects the stability of the training. To mitigate this issue, we normalize it by dividing its norm:
$\nabla_{\boldsymbol{a}_{t}}Q_{\text{norm}}(\boldsymbol{s},\boldsymbol{a}_{t})=\frac{\nabla_{\boldsymbol{a}_{t}}Q(\boldsymbol{s},\boldsymbol{a}_{t})}{||\nabla_{\boldsymbol{a}_{t}}Q(\boldsymbol{s},\boldsymbol{a}_{t})||+\epsilon},$ (14)
where $\epsilon$ is a small constant to prevent division by zero.
Ultimately, we construct the Q-gradient field objective function to facilitate the training of the diffusion policy, where $\pi_{\theta}(\boldsymbol{a}_{t}|\boldsymbol{s})$ denotes the action generated using the diffusion policy as defined in Eq. (6):
$\mathcal{L}_{g}(\theta)=\min_{\theta}\underset{\begin{subarray}{c}\boldsymbol{s}\sim\mathcal{B}\\ t\sim\text{U}(1,T)\\ \boldsymbol{a}_{t}\sim\pi_{\theta}(\boldsymbol{a}_{t}|\boldsymbol{s})\end{subarray}}{\mathbb{E}}\left[\left\|w(t)\nabla_{\boldsymbol{a}_{t}}Q_{\text{norm}}(\boldsymbol{s},\boldsymbol{a}_{t})-S_{\theta}(\boldsymbol{s},\boldsymbol{a}_{t},t)\right\|_{2}^{2}\right],$ (15)
3.3 DACER2: A High Efficiency Diffusion RL Algorithm
In the critic component, we adopt the double Q-learning strategy fujimoto2018addressing to alleviate overestimation bias. Specifically, we maintain two independent Q-function estimators, denoted as $Q_{\phi_{1}}(\boldsymbol{s},\boldsymbol{a})$ and $Q_{\phi_{2}}(\boldsymbol{s},\boldsymbol{a})$, which are trained to approximate the true action-value function. To enhance training stability, we introduce two corresponding target networks, $Q_{\bar{\phi}_{1}}(\boldsymbol{s},\boldsymbol{a})$ and $Q_{\bar{\phi}_{2}}(\boldsymbol{s},\boldsymbol{a})$, which are updated softly from the main networks following the technique in van2016deep .
The Q-networks are optimized by minimizing the Bellman error. For each network $Q_{\phi_{i}}(\boldsymbol{s},\boldsymbol{a})$, the loss $J_{Q}(\phi_{i})$ is defined as:
$J_{Q}(\phi_{i})\!=\!\underset{\begin{subarray}{c}(\boldsymbol{s},\boldsymbol{a},\boldsymbol{r},\boldsymbol{s}^{\prime})\sim\mathcal{B}\\ \boldsymbol{a}^{\prime}\sim\pi_{\theta}(\boldsymbol{a}_{0}|\boldsymbol{s})\end{subarray}}{\mathbb{E}}\left[\!\left(\!\left(r(\boldsymbol{s},\boldsymbol{a})\!+\!\gamma\min_{i=1,2}Q_{\bar{\phi}_{i}}(\boldsymbol{s}^{\prime},\boldsymbol{a}^{\prime})\!\right)-\!Q_{\phi_{i}}(\boldsymbol{s},\boldsymbol{a})\!\right)^{2}\right],\$ (16)
where $\gamma$ is discount factor, the target is computed as the smaller of the two target Q-values, $Q_{\bar{\phi}_{1}}(\boldsymbol{s}^{\prime},\boldsymbol{a}^{\prime})$ and $Q_{\bar{\phi}_{2}}(\boldsymbol{s}^{\prime},\boldsymbol{a}^{\prime})$, to prevent over-optimistic estimates. Furthermore, we incorporate the distributional value estimation framework from DSAC duan2023dsac to further mitigate overestimation issues.
[…]
where $\eta$ is a hyperparameter, $\mathcal{L}_{q}(\theta)=\mathbb{E}_{\boldsymbol{s}\sim\mathcal{B},\boldsymbol{a}_{0}\sim\pi_{\theta}(\cdot|\boldsymbol{s})}\left[-Q_{\phi}(\boldsymbol{s},\boldsymbol{a}_{0})\right],$ $p(s)$ is a distribution over states.
4 Experimental Results
This section presents the experimental results. We first evaluate the multimodality of DACER2 and DACER using a toy environment called "Multi-goal" haarnoja2017reinforcement , as illustrated in Fig. 2. Then we show the empirical results of the proposed DACER2 algorithm evaluated with OpenAI Gym MuJoCo 2012MuJoCo tasks. With these experimental results, we aim to answer three questions:
  • Does DACER2 demonstrate stronger multimodal capabilities compared to DACER?
  • How does DACER2 compare to previous popular online RL algorithms and existing diffusion-based online RL algorithms?
  • How does the inference and training efficiency of DACER2 compare with existing diffusion-based RL methods?
Baselines.
The baselines encompass two categories of model-free reinforcement learning algorithms. The first category consists of diffusion-based RL methods, including a range of recent diffusion-policy online algorithms such as DACER wang2024diffusion , QVPO ding2024diffusion , DIPO yang2023diffusion , and QSM psenka2023learning . The second category includes classic model-free online RL baselines, namely TD3 fujimoto2018addressing , SAC haarnoja2018soft , PPO schulman2017proximal , and DSAC duan2023dsac .
Evaluation Setups.
We implemented our algorithm in JAX and evaluated it on eight MuJoCo tasks using the same metrics as DACER. Experiments were conducted on a system equipped with an AMD Ryzen Threadripper 3960X 24-core processor and an NVIDIA GeForce RTX 4090 GPU. For example, training Walker2d for 1.5M steps takes about two hours. Performance was measured by averaging the top returns from the final 10% of iterations over ten episodes. Results across five random seeds are reported with mean and standard deviation. For classic model-free baselines, we cite DACER-reported results, while all diffusion-based methods were re-evaluated. Furthermore, the training curves presented in Fig. 3 demonstrate the stability of the training process.
4.1 Toy Example
We evaluate the trained policy in the "Multi-goal" environment by initializing the agent at the origin and sampling 100 trajectories. The agent is tasked with navigating toward one of six symmetrically arranged goal points. As illustrated in Fig. 2, DACER2 demonstrates significantly enhanced multimodality compared to DACER. Specifically, DACER2 consistently reaches all six target locations with approximately uniform coverage, indicating that it effectively captures the underlying multimodal structure of the task and maintains balanced expression across distinct behavioral modes.
In contrast, DACER only reaches five out of the six goals, and its visitation frequencies are noticeably uneven. This imbalance suggests mode collapse and reveals limitations in its ability to represent diverse behaviors. These results highlight that DACER2 is better equipped to express diverse, mode-separated policies in multimodal environments. Additional multimodal experiments are provided in Appendix C.
(a) DACER2
(b) DACER
Figure 2: Multi-goal Task. Trajectories generated by a policy learned using our method (solid blue lines) are shown, with the $x$-axis and $y$-axis representing 2D positions (states). The agent is initialized at the origin, and the goals are marked as red dots. The level curves indicate the reward, and reaching within 1 of the endpoint signifies task completion. On the left are the experimental results of DACER2, and on the right are the experimental results of DACER.
4.2 Comparative Evaluation
The performance comparison, as summarized in Table 1 and illustrated in Fig. 3, indicates that our method consistently outperforms all baseline algorithms across the OpenAI Gym MuJoCo benchmark tasks. In particular, in more challenging locomotion environments such as Humanoid, Ant, HalfCheetah, and Walker2d, our method achieves substantial improvements over SAC, with respective gains of 27.4%, 43.4%, 9.3%, and 24.1%.
Furthermore, among diffusion-based online RL approaches, DACER2 further improves upon DACER, achieving performance gains of 0.2%, 15.3%, 3.7%, and 16.4% in the same set of tasks. Notably, DACER2 attains performance that is comparable to or even surpasses DACER, while requiring only five diffusion steps. These results collectively demonstrate the robustness and efficiency of our method, and further underscore the potential of diffusion-based policies for online RL.
(a) Humanoid-v3
(b) Ant-v3
(c) HalfCheetah-v3
(d) Walker2d-v3
(e) Inverted2Pendulum-v3
(f) Hopper-v3
(g) Pusher-v2
(h) Swimmer-v3
Table 1: Average final return. Computed as the mean of the highest return values observed in the final 10% of iteration steps per run. The value for each task is bolded. ± corresponds to standard deviation over five runs.
Humanoid Ant HalfCheetah Walker2d
Classic Model-Free RL PPO $6869\pm 1563$ $6157\pm 185$ $5789\pm 2201$ $4832\pm 638$
TD3 $5632\pm 436$ $6184\pm 487$ $8633\pm 4041$ $5237\pm 336$
SAC $9335\pm 696$ $6427\pm 804$ $16573\pm 224$ $6201\pm 263$
DSAC $10829\pm 243$ $7086\pm 261$ $17025\pm 157$ $6424\pm 147$
Diffusion Policy RL QSM $6072\pm 691$ $4783\pm 1235$ $11401\pm 882$ $5685\pm 437$
DIPO $9353\pm 356$ $3449\pm 149$ $12267\pm 2180$ $5066\pm 365$
DACER $11868\pm 56$ $7994\pm 61$ $17466\pm 96$ $6610\pm 12$
QVPO $9656\pm 252$ $6484\pm 145$ $14355\pm 175$ $6057\pm 352$
DACER2 (ours) $\mathbf{11896\pm 53}$ $\mathbf{9217\pm 64}$ $\mathbf{18107\pm 146}$ $\mathbf{7694\pm 352}$
IDP Hopper Pusher Swimmer
Classic Model-Free RL PPO $9357\pm 2$ $2647\pm 481$ $-22.9\pm 1.4$ $130\pm 2$
TD3 $9347\pm 15$ $3569\pm 455$ $-21.4\pm 1.2$ $134\pm 5$
SAC $\mathbf{9360\pm 0}$ $2483\pm 943$ $-19.6\pm 0.3$ $140\pm 14$
DSAC $\mathbf{9360\pm 0}$ $3660\pm 533$ $\mathbf{-19.4\pm 0.9}$ $138\pm 6$
Diffusion Policy RL QSM $591\pm 98$ $2006\pm 251$ $-73.8\pm 3.0$ $46\pm 1$
DIPO $9355\pm 2$ $3839\pm 40$ $-20.7\pm 0.2$ $55\pm 2$
DACER $9354\pm 2$ $4094\pm 54$ $-19.7\pm 0.2$ $150\pm 4$
QVPO $9354\pm 5$ $4046\pm 94$ $-20.4\pm 0.1$ $130\pm 10$
DACER2 (ours) $9359\pm 1$ $\mathbf{4213\pm 91}$ $\mathbf{-19.4\pm 0.1}$ $\mathbf{165\pm 2}$
4.3 Efficiency Analysis
We evaluate the training and inference efficiency of the algorithm, as illustrated in Fig. 4. All algorithms are implemented and tested using PyTorch to reflect practical deployment conditions, as PyTorch is commonly adopted in real-world control systems for both training and inference. Compared to the three diffusion-based algorithms, DACER2 is 41.7%, 49.3%, and 50.0% faster than DACER, QVPO, and DIPO, respectively, but 16.7% slower than QSM in terms of training speed. Given that the performance of our algorithm significantly outperforms QSM, the minor difference in training speed can be disregarded. Regarding inference speed, DACER2 shows a 60.6%, 82.5%, 60.6%, and 60.6% improvement over DACER, QVPO, QSM, and DIPO, respectively.
DACER2 adopts a Q-gradient field to assist the score function training under the max-Q objective framework, allowing more effective guidance during each denoising step. This design improves sample efficiency and accelerates convergence without requiring a large number of diffusion steps. As a result, even with only five diffusion steps, our algorithm achieves strong performance and substantial gains in training and inference efficiency. In real-time industrial control tasks, the inference time should be less than 1 millisecond to meet control requirements. Among the existing diffusion-based algorithms, only DACER2 meets this constraint.
Figure 4: Comparison of inference and training time. The left subfigure depicts the time required for each algorithm to complete a training step, excluding the time spent interacting with the environment. The right subfigure depicts the time required for the policy network to output an action after receiving a single state as input. All results are based on five different seeds.
4.4 Ablation Study
In this section, we conduct ablation study to investigate the impact of the following two aspects on the performance of the diffusion policy: 1) whether to use the Q-gradient field training objective function derived from the connection between noise prediction network and Q-gradient field; 2) whether to weight $\nabla_{a}Q(s,a)$ according to the time step of the diffusion process.
(a) Ablation for the Q-gradient field training objective function.
(b) Ablation for the time-weighted $\nabla_{a}Q(s,a)$.
Figure 5: Ablation experiment curves. (a) DACER2 represents the algorithm with Q-gradient field loss, and 20 steps represents the diffusion step size of 20. DACER2’s performance on Walker2d-v3 is far better than DACER. (b) Time-weighted means using $w(t)\nabla_{a}Q(s,a)$ instead of $\nabla_{a}Q(s,a)$ as the target value in the Q-gradient field training loss.
Q-gradient field training objective function.
In this ablation study, we first fixed the diffusion step size at 5 to evaluate the impact of the Q-gradient field loss function. As shown in Fig. 5(a), removing this objective—yielding a variant equivalent to DACER with 5 steps—led to a significant decline in performance. Next, we increased the diffusion step size of the algorithm without the Q-gradient field loss to 20; however, its performance still remained inferior to that of DACER2 with only 5 diffusion steps. These results demonstrate that the Q-gradient field loss function plays a crucial role in guiding the diffusion denoising process and serves as a key component for enhancing overall performance.
Time-weighted $\nabla_{a}Q(s,a)$.
We conducted an experiment to demonstrate that using time-weighted $\nabla_{a}Q(s,a)$ can further improve performance. As shown in Fig. 5(b), directly using $\nabla_{a}Q(s,a)$ as the target value in the Q-gradient field training loss, instead of the $w(t)\nabla_{a}Q(s,a)$, results in performance degradation. This is because different timesteps require matching different magnitudes of noise prediction, which enhances both training stability and final performance.
5 Related Work
In this section, we review existing works on using the diffusion model as a policy function in combination with RL.
Online RL with Diffusion Policy.
Online RL enables agents to refine their policies through real-time interaction. Yang et al. introduced DIPO yang2023policy , which maintains a dedicated diffusion buffer to store actions and model them using diffusion techniques. Psenka et al. proposed QSM psenka2023learning , which aligns policies with $\nabla_{a}Q$ via score matching, but is sensitive to value gradient inaccuracies across the action space. Recently, Ding et al. ding2024diffusion proposed QVPO, which weights diffusion-sampled actions by Q-values without computing gradients. However, it uses a fixed ratio of uniform samples to boost the entropy, lacking adaptive control and later degrading performance. Ma et al. ma2025soft proposed SDAC, which uses score matching over noisy energy-based diffusion. It avoids requiring optimal actions but suffers from high gradient variance due to poor sampling in high-Q regions.
Offline RL with Diffusion Policy.
Offline RL focuses on learning optimal policies from suboptimal datasets, with the core challenge being the out-of-distribution (OOD) problem kumar2020conservative ; fujimoto2019off . Diffusion models are naturally suited for offline RL due to their ability to model complex data distributions. Wang et al. proposed Diffusion-QL wang2022diffusion , which combines behavior cloning through a diffusion loss with Q-learning to improve policy learning. However, Diffusion-QL suffers from slow training and instability in OOD regions. To address the former, Kang et al. proposed Efficient Diffusion Policy (EDP) kang2023efficient , which speeds up training by initializing from dataset actions and adopting a one-step sampling strategy. To mitigate OOD issues, Ada et al. introduced SRDP ada2024diffusion , which enhances generalization by integrating state reconstruction into the diffusion policy. Furthermore, Chen et al. proposed CPQL chen2023boosting , a consistency-based method that improves efficiency via one-step noise-to-action generation during both training and inference, albeit with some performance trade-offs.
Comparison with DACER.
Wang et al. proposed DACER wang2024diffusion , which uses the reverse diffusion process as a policy approximator and employs a Gaussian mixture model (GMM) to estimate policy entropy for balancing exploration and exploitation. However, the work lacks theoretical justification for why maximizing the expected Q-value enables learning multimodal policies. Moreover, a key trade-off remains: long diffusion processes hinder training efficiency, while fewer steps lead to performance degradation. Our work addresses this by reducing the diffusion steps while maintaining or even improving both performance and policy multimodality.
6 Conclusion
In this paper, we address the critical challenge of balancing inference efficiency and performance in diffusion-based online RL. By introducing a Q-gradient field objective and a time-dependent weighting scheme, our method enables each denoising step to be guided by the Q-function with adaptive emphasis over time. This design allows the policy to achieve strong performance with only five diffusion steps, significantly improving both training and inference speed. In addition, our framework inspires a novel class of SDE-based policy functions that maximize the expected Q-value under a fixed entropy constraint, enabling effective learning of expressive soft policy distributions.
Appendix A Theoretical Results
Proof. Let $\mathcal{S}$ be the state space, $\mathcal{A}$ denotes the continuous action space, and $p(s)$ is a probability measure over states. We seek a family of policies $\{\pi(\cdot\mid s)\}_{s\in\mathcal{S}}$ belonging to the constrained space:
@@ Appendix A Theoretical Results
[…]
@@
[…]
@@
[…]
@@
Appendix B Environmental Details
MuJoCo [31]:

### v2

@@ Appendix A Theoretical analysis
Proof. We seek a family of policies $\{\pi(\cdot\mid s)\}_{s\in\mathcal{S}}$ belonging to the constrained space:
@@
$\mathcal{L}\bigl(\{\pi(\cdot\mid s)\},\alpha,\{\eta(s)\}\bigr)$ $=\int_{\mathcal{S}}\!\int_{\mathcal{A}}\Bigl[p(s)\pi(a\mid s)Q(s,a)-\alpha\,p(s)\pi(a\mid s)\log\pi(a\mid s)+\eta(s)\pi(a\mid s)\Bigr]\,da\,ds$ (20)
$\quad-\alpha\,\mathcal{H}^{\text{global}}_{0}-\int_{\mathcal{S}}\eta(s)\,ds.$
@@
Imposing $\int_{\mathcal{A}}\pi(a\mid s)\,da=1$, we can determine
$C(s)=\Bigl[\int_{\mathcal{A}}\!\exp\!\bigl(Q(s,a^{\prime})/\alpha\bigr)\,da^{\prime}\Bigr]^{-1}.$ (25)
@@
$\pi^{*}(a\mid s)=\frac{\exp\!\bigl(Q(s,a)/\alpha\bigr)}{\displaystyle\int_{\mathcal{A}}\!\exp\!\bigl(Q(s,a^{\prime})/\alpha\bigr)\,da^{\prime}}\qquad\forall s\in\mathcal{S},\;a\in\mathcal{A}.$ (26)
@@
Appendix B Related Work
We review existing works on using the diffusion model as a policy function in combination with RL.
Online RL with Diffusion Policy.
Online RL enables agents to refine their policies through real-time interaction. Yang et al. introduced DIPO (Yang et al., 2023a), which maintains a dedicated diffusion buffer to store actions and model them using diffusion techniques. Psenka et al. proposed QSM (Psenka et al., 2023), which aligns policies with $\nabla_{a}Q$ via score matching, but is sensitive to value gradient inaccuracies across the action space. Recently, Ding et al. (Ding et al., 2024) proposed QVPO, which weights diffusion-sampled actions by Q-values without computing gradients. However, it uses a fixed ratio of uniform samples to boost the entropy, lacking adaptive control and later degrading performance. Ma et al. (Ma et al., 2025) proposed SDAC, which uses score matching over noisy energy-based diffusion. It avoids requiring optimal actions but suffers from high gradient variance due to poor sampling in high-Q regions. Celik et al. proposed DIME (Celik et al., 2025), which derives a lower bound on the diffusion policy entropy and integrates it into the maximum-entropy RL framework. However, directly incorporating an inaccurate entropy estimate into the policy objective can degrade performance.
Offline RL with Diffusion Policy.
Offline RL focuses on learning optimal policies from suboptimal datasets, with the core challenge being the out-of-distribution (OOD) problem (Kumar et al., 2020; Fujimoto et al., 2019). Diffusion models are naturally suited for offline RL due to their ability to model complex data distributions. Wang et al. proposed Diffusion-QL (Wang et al., 2023), which combines behavior cloning through a diffusion loss with Q-learning to improve policy learning. However, Diffusion-QL suffers from slow training and instability in OOD regions. To address the former, Kang et al. proposed Efficient Diffusion Policy (EDP) (Kang et al., 2023), which speeds up training by initializing from dataset actions and adopting a one-step sampling strategy. To mitigate OOD issues, Ada et al. introduced SRDP (Ada et al., 2024), which enhances generalization by integrating state reconstruction into the diffusion policy. Furthermore, Chen et al. proposed CPQL (Chen et al., 2023), a consistency-based method that improves efficiency via one-step noise-to-action generation during both training and inference, albeit with some performance trade-offs.
Comparison with DACER.
Wang et al. proposed DACER (Wang et al., 2024), which uses the reverse diffusion process as a policy approximator and employs a GMM to estimate policy entropy for balancing exploration and exploitation. However, the work lacks a theoretical justification for why maximizing the expected Q-value enables learning multimodal policies. Moreover, a key trade-off remains: long diffusion processes hinder training efficiency, while fewer steps lead to performance degradation. Our work addresses this by reducing the diffusion steps while maintaining or even improving both performance and policy multimodality.
Comparison with QSM.
Psenka et al. proposed QSM (Psenka et al., 2023), an algorithm that aligns diffusion model policies with $\nabla_{\bm{a}}Q(\bm{s},\bm{a})$ by leveraging their score-based structure. Both methods leverage Q-gradients for diffusion policy optimization. QSM employs score matching, whereas DACERv2 performs end-to-end Q-value maximization augmented with a time-weighted score-matching loss and entropy regularization, resulting in a multi-task objective. DACERv2 additionally stabilizes Q-gradients through normalization and improves efficiency, converging in just 5 diffusion steps compared to QSM’s approximately 20.
Appendix C Environmental Details
MuJoCo (Brockman et al., 2016):